A coffee cup calorimeter is prepared, containing 100.000 g of water (specific heat capacity = 4.184 J/g K) at initial temperature 80.000 C. A salt weighing 5.398 g is quickly added. The salt has a molar mass of 497.886 g/mol. The final temperature of the solution is 8.675 C.

Assume no heat loss to the surroundings. Assume the specific heat capacity of the solution is equal to that of pure water, and that the mass of the solution is equal to the mass of the solid plus the mass of water in the calorimeter.

What is the molar heat of solution for the salt, in kJ/mol?

Report your answer to three digits after the decimal.
Answer #1

Salt - Water mars (ml=5.398 mars (m/= 100gm. gm Specific Heat Capauty ($1= 4:1845[gm-k i nitial temp (Til = 80°C final temp (H₂O is equal to heat of solution ( AH sow) (AH) sono = Heat lose by Soln (H₂O) on adding Salt (AH) go w ² MSAT m = mars of soS = Specific heat capacity = 4:184.J]m-ke AT = temperature difference = 80-8.675 = 71.325 (OH)8qw = 105.398x4•184x71325 = 31,We have 5.398 gm of salt 80, Mots of salt bivermars - molar mars = 5.398 497.886 moles = 0.010842 So, on addition of 0.010842= 2901076 KI So, Enthalpy of solution 7 orgoma= 4901.076 km mo 1